Four time the smallest of three consecutive odd integers is 236 more than the sum of the other two intergers. Find the integers.
Answer by checkley77(12844)

Let x, x+2 & x+4 be the 3 odd integers.
4x=236+(x+2)+(x+4)
4x=236+2x+6
4x-2x=242
2x=242
x=242/2
x=121 ans. for the smallest integer.
121+2=123 for the middle integer.
121+4=125 for the largest integer.
Proof;
4*121=236+123+125
484=484